Solve the system by substitution.
x + y = 12
y = 2x

Answer by malakumar_kos@yahoo.com(315)


Solve the system by substitution.
x + y = 12
y = 2x

x+y = 12.......(1)
y = 2x.......(2)
substituting fo y in eq'n (1) we get x+2x = 12
3x = 12
x = 12/3 = 4
substituting for x in eq'n (2) we get y = 2(4) = 8
The values of x = 4 and y = 8 are the required solutions for the eq'ns